Several people were out tonight: Nathan, Ankit, Laura, Michael and I.
There were camera connection problems early in the evening that got resolved around 11:30PM, followed by TCS freezing up. We went through a manual shutdown/re-zenithing of the telescope. Pointing was a bit off so we got the location of the telescope by taking a 30 second exposure that we uploaded to astrometry.net to get a RA/Dec.
After that the evening was smooth for a while. Observed a maximum of EY UMa, then some M82, but in both cases clouds started to get in the way.
The standard field SA104 was also observed.
Then the filter wheel froze up…and spontaneously started working again about 15 minutes later.
Took diffuser flats in V, B, R, then left around 6:30 AM.